OpenStudy (anonymous):

simplify sin Θ/√1-sin Θ

OpenStudy (anonymous):

multiply the top and bottom by \(\sqrt{1-\sin\theta}\).

OpenStudy (anonymous):

Is it tan^2 Θ ?

OpenStudy (anonymous):

Hmmm, actually maybe do \(\sqrt{1+\sin\theta}\)

OpenStudy (anonymous):

\[=\frac{ 2\sin \frac{ \theta }{ 2 }\cos \frac{ \theta }{ 2 } }{ \sqrt{\cos ^2 \frac{ \theta }{ 2 }+\sin ^2\frac{ \theta }{ 2 }-2\sin \frac{ \theta }{ 2 }\cos \frac{ \theta }{ 2 }} }\] \[=\frac{ 2 \sin \frac{ \theta }{ 2 }\cos \frac{ \theta }{ 2} }{ \left( \cos \frac{ \theta }{2 }-\sin \frac{ \theta }{2 } \right) }\]

OpenStudy (anonymous):

\[ \frac{\sin \theta \sqrt{1+\sin\theta}}{\sqrt{1-\sin^2\theta}} = \frac{\sin \theta \sqrt{1+\sin\theta}}{|\cos\theta|} \]

OpenStudy (anonymous):

tan Θ tan2 Θ 1 −1 These are the choices.

OpenStudy (anonymous):

there is something wrong in your statement.

OpenStudy (anonymous):

You're right. I just checked. The 1-sin Θ is actually 1-sin^2Θ I apoligize, It didn't paste correctly.

OpenStudy (anonymous):

is it \[\frac{ \sin \theta }{ \sqrt{1-\sin ^2\theta } }?\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

very simple use \[\sin ^2\theta+\cos ^2\theta=1\]

OpenStudy (anonymous):

Wow, that was simple. Thank you so much for all of your time.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[opp^2+adj^2=hyp^2 \\ \text{ and we know the } adj \text{ and the hyp } \\ \text{ just solve for opp}\]